# Comparing the latent features of universal machine-learning interatomic potentials

### v1

@@ III.2 Comparing model variants
Beyond the comparison between entirely different uMLIPs, we assess the consistency of information content across uMLIP “variants” of the same architecture. By computing the reconstruction errors between different instances of the same architecture, we establish the sensitivity of various architectures to changes in the training dataset or targets. We also establish a scale for the “internal” variability of each architecture that helps calibrate the comparison between models that differ more substantially.
@@
We categorize the variants into: “single-task” models where a separate model architecture is trained from scratch for distinct datasets, and “multi-task” models that either adopt a multi-head Caruana1997; ruder2017 or mixture-of-experts jacobs1991adaptive; riquelme2021scaling; shazeer2017outrageously approach to simultaneously target multiple datasets with a single model architecture, often with significant weight sharing within the model architecture between the prediction tasks.
@@
Figure 2: Reconstruction errors among single-task MACE models trained on MPtrj (MACE-MP-0b3), OMat24, and MatPES (PBE and r2SCAN) datasets, evaluated on the MAD test MPTRAJ; MATPES; omat24.
@@
We first compare the reconstruction errors across the single-task MACE variants, which have been trained separately on MPtrj MPTRAJ, OMat24 omat24, MatPES-PBE, and MatPES-r2SCAN MATPES. The latter two datasets contain the same structures, but the reference calculations have been performed with PBE perd+96prl and r2SCAN Furness2020 functionals, respectively. Overall, the reconstruction errors are moderately low across the variants (Fig. 2) (the average off-diagonal GFRE is 0.31, and the average LFRE is 0.12). The MPtrj variant exhibits the largest overall GFREs and LFREs (average GFRE and LFRE are 0.42 and 0.15, respectively), which reflects the limited information content of the smallest and oldest MPtrj dataset. The reconstruction errors between MatPES-PBE and MatPES-r2SCAN are comparable to the other cases, revealing that having different functionals in the reference calculations converges the resulting variant models into significantly distinct latent spaces. The lowest overall errors are observed for the OMat24 variant (on average, the GFRE is 0.25 and the LFRE is 0.09), which is indicative of the relatively rich information content of OMat24 dataset that is often utilized in the pre-training of uMLIPs eSEN; GRACE; mace_multihead; ORB3; omat24.
@@
In DPA-3.1, a model trained on the OpenLAM-v1 collection of 31 datasets OpenLAMv1, the following multi-task learning strategy is adopted: the model incorporates a shared backbone network while still being conditioned for different datasets via one-hot encoding. This backbone is followed by multiple specialized output heads, or branches, tailored to each dataset DPA3. Reconstruction results between branches corresponding to OMat24, MPtrj, OC20, ODAC, and SPICE datasets omat24; MPTRAJ; OC20; ODAC; SPICE, presented in SI VII.1.1, indicate comparable latent space consistency to the MACE single-task models, suggesting that strong parameter sharing without dynamic routing preserves highly universal latent representations.
@@
UMA is a uMLIP developed with the goal of having a single model that performs well across all atomistic modeling tasks, encompassing materials, molecules, and their interactions. To do so, the model has been trained simultaneously on OMol25, OMat24, OC20, ODAC25, and OMC25 OMOL; omat24; OC20; ODAC; OMC25. Instead of adopting a multi-head approach for the different datasets, the model incorporates MoLE strategy in its interaction layers, where the weights of the different experts are computed based on the global embedding that encodes the dataset-specific information UMA. UMA-S-1P1 shows greater variability in reconstruction errors (see Fig. 3), which suggests that the MoLE mechanism encourages stronger specialization of feature manifolds corresponding to each dataset. We observe that the OMat24 task reconstructs other variants with relatively low errors (average off-diagonal GFRE is 0.49 and LFRE is 0.34), similar to previous results with MACE, whereas other variants struggle to approximate its latent space linearly. We also observe group-level consistency, as the features of catalysis-related tasks (OC20, ODAC25, OMC25) generally exhibit lower reconstruction errors (GFRE is 0.46, and LFRE is 0.34 on average across this group).
@@
Figure 3: Reconstruction errors among different input tasks of UMA-S-1P1 trained on OMat24, OMOL, OC20, ODAC, and OMC omat24; OMOL; OC20; ODAC; OMC25, evaluated on the MAD test.
@@
To investigate the influence of the training domain on the learned latent space, we compared the organic-focused MACE-OFF23 model MACEOFF with the materials-focused MACE-MP-03b MACE by evaluating the reconstruction errors on the organic structures of the MAD test subset. As detailed in SI VII.2 (Table S2), the materials-trained model reconstructs the organic-trained model almost perfectly (GFRE is 0.02), whereas the reverse direction yields remarkably higher global error (GFRE is 0.36), despite similar local agreement (LFRE is 0.04). This could suggest that the broader materials manifold learned by MACE-MP-03b cannot be fully captured by MACE-OFF23, which has been exposed only to organic systems.
@@
We further examined the effect of model capacity by comparing the MACE checkpoints of varying sizes (small, medium, large) within the same architecture family, for both the materials-focused MACE-MP-0a series and the organic-focused MACE-OFF23 series (SI VII.3, Fig. S5). Interestingly, regardless of size, all MACE-MP-0a variants reconstruct each other with near-perfect linear mapping, whereas reconstructing the large MACE-OFF23 checkpoint from its smaller counterparts is significantly harder (average GFRE is 0.48 when small/medium $\rightarrow$ large). A similar size-scaling analysis on PET-OMAT models Bigi2025prep (SI Fig. S6) shows that smaller models can still approximate the principal manifold of larger ones reasonably well, with the medium-sized variant offering a good information content/size tradeoff.
@@
Finally, we assessed the impact of single-task versus multi-head readout architectures using two MACE models trained on the same OMat24 dataset: the original single-head MACE-OMAT-0 and the OMat-specific head of the multi-head MACE-MH-1 model mace_multihead (the latter having undergone multi-head replay fine-tuning after initial pre-training). Results are presented in SI VII.4 (Table S4). Both directions exhibit moderate global reconstruction error (GFRE is $\sim$ 0.33) but low local error (LFRE is less than 0.10), demonstrating that the multi-head replay protocol preserves the core local feature neighborhoods of the original single-head representations while allowing the readout to become more expressive.
@@ III.3 Varying targets
Although we have primarily focused on uMLIPs, universal models are also available for properties other than the ground state potential energy surface. To this end, we compare the latent features of PET-MAD and PET-MAD-DOS PET-MAD-DOS-2025. The two models have nearly identical architectures, differing only in the last linear readout layer due to the different sizes of the targets (last layer features are still of the same dimension), and are both trained on the MAD dataset. While PET-MAD was trained on energies and forces, PET-MAD-DOS was trained on the electronic density of states (DOS).
The results in SI VIII reveal that PET-MAD-DOS exhibits more information content than PET-MAD, as seen by the lower GFRE and LFRE values (0.68 and 0.39 for energy $\to$ DOS vs. 0.56 and 0.28 for DOS $\to$ energy). While this observation is expected, provided that DOS contains far more information than the energies and forces, it is still interesting to see that the information content in the latent features also reflects our physical knowledge of the targets. Still, the substantial reconstruction error in the DOS $\to$ energy direction indicates that the energy-based latent features also encode distinct information.
@@ III.4 Fine-tuning
The advent of uMLIPs has made fine-tuning on domain-specific datasets a standard approach for obtaining accurate potentials tailored to specific applications that can still generalize well deng2024; Jacobs2025; Radova2025; Kaur2025. Here, we analyze the reconstruction errors between features that arise from different fine-tuning strategies, as well as a model trained entirely from scratch. We also study the evolution of learned latent features along a “training trajectory” by computing the reconstruction errors of latent features sampled from model checkpoints at intermediate training epochs with respect to the fully trained model.
@@
For our analysis, we consider the PET architecture and its pre-trained PET-MAD uMLIP as the baseline. We train material-specific MLIPs for lithium thiophosphate (LPS), a promising class of solid-state electrolytes, widely studied for the next generation of battery materials liuAnomalousHighIonic2013a; fragapane2025lipselectrolytematerialsbenchmark; katoLithiumionconductiveSulfidePolymer2021; manthiramLithiumBatteryChemistries2017; tisiThermalConductivityLi2024; gigliMechanismChargeTransport2024. We consider the following training or fine-tuning procedures: a bespoke model trained “from scratch” on the dataset from randomly initialized weights; full fine-tuning (FF), where all trainable model weights are optimized on the new dataset; “head-only” fine-tuning (HF), where only the weights of the MLP head are optimized during training while the rest remains fixed; full transfer learning (FTL), where a new MLP head is added with random weight initialization, and then all trainable weights of the model are optimized; head transfer learning (HTL), where a new MLP head is created with random weights, and only the MLP head weights are optimized. A schematic overview of the trainable parts for these fine-tuning strategies is given in Fig. 4. The training procedures were carried out using metatrain metatrain (see SI IX for final errors). We use the dataset from Ref. tuerk2025, which was previously used to model complex surface structural reconstructions on LPS. The dataset contains 4,088 structures, yielding 225,905 atomic environments in total. In particular, it contains the structures of bulk LPS in each of its phases ($\gamma$, $\beta$, $\alpha$, and amorphous) and the structures of the LPS surface with and without absorbed water molecules.
@@
Figure 4: Schematic overview of fine-tuning strategies for the PET model architecture: full fine-tuning, head fine-tuning, full transfer learning, and head transfer learning. GNN based on the Point Edge Transformer (PET) computes the backbone (BB) features (general, transferable node and edge representations). For a given prediction target, these BB features are fed into the readout head (MLPs) to generate the last-layer (LL) features (target-specific node and edge representations). The four illustrated strategies involve training different combinations of these components. The trainable parts corresponding to each strategy are colored.
@@
Fig. 5 shows the reconstruction errors among the final models from the described fine-tuning strategies, the pre-trained PET-MAD model, and a bespoke PET model trained from scratch on the LPS dataset. The reconstruction errors are notably low between all of the fine-tuned models (FF, HF, FTL, and HTL) and the pre-trained PET-MAD model, which signifies a high degree of similarity in their last-layer features despite differences in training protocols. This suggests that fine-tuning retains a strong bias from the pre-trained model. The bespoke model exhibits the highest reconstruction errors, yet the values are still moderate, revealing that its training has converged to a distinct yet nearby minima with respect to the other models.
@@
Figure 5: Reconstruction errors (GFRE and LFRE) of last-layer atom-level features between PET checkpoints on the LPS dataset.
@@
Fig. 6 shows the evolution of the reconstruction errors along the training trajectory computed relative to the fully converged model for each strategy. The bespoke model trained from scratch starts with high GFRE and LFRE, which gradually decay, yet remain above 0.09 for GFRE and 0.05 for LFRE even after 1000 epochs. In contrast, fine-tuned strategies begin with lower initial reconstruction errors due to pre-conditioning with the PET-MAD uMLIP, which then converge quickly to near-zero values. In FTL and HTL, the curve is not monotonic, and more than 400 epochs are needed for convergence, most likely due to the randomized head initialization. In HF, a similar number of epochs is needed to align the features with those of the final model, whereas FF converges the reconstruction errors to <0.01<0.01 with only 20 epochs. This can be explained by the relative differences in the training degrees of freedom between the two strategies. Overall, the combination of low reconstruction errors in Fig. 5 and rapid convergence in Fig. 6 demonstrates that the last-layer features of PET-MAD provide a robust starting point for the finetuning strategies. This explains why uMLIPs converge rapidly to high accuracy when fine-tuned to a new, narrower dataset Kaur2025.
@@
Figure 6: Evolution of the reconstruction errors of last-layer atom-level features during training on the LPS dataset, relative to the fully trained model of each fine-tuning strategy.
@@
While we have chosen the last-layer (LL) features for our analysis thus far, the latent features can be extracted from any other layer within the NN architectures of the uMLIPs. Another features of particular interest are the “backbone” (BB) latent features, which correspond to the representations obtained immediately after all message-passing iterations, but before entering a multilayer perceptron (MLP) head, which is an architectural choice followed by many uMLIPs (e.g., see Fig. 4 for a schematic representation of PET-MAD). The BB features are of particular interest in multi-head and multi-task models, where they serve as a shared input across the multiple heads. To provide examples of both single-task and multi-task contexts, we evaluate the reconstruction errors between the last-layer features and the backbone features for PET and UMA.
@@
For PET (Fig. S8 in SI X), both reconstruction errors are higher and more varied across model variants compared to UMA-S-1P1, indicating greater divergence between BB and LL representations. Even in the GFRE, reconstructing BB from LL yields higher errors than the reverse case. Once again, the LFRE is lower in both reconstruction tasks, but the improvement is only minor in the case of LL $\rightarrow$ BB. We speculate that the large information loss might be associated with the MLP eliminating the non-invariant parts of the BB features, which takes place as the PET model learns the symmetries of the energy prediction task.
@@
Figure 7: Global and local reconstruction errors for last-layer features ($L$) and backbone features ($B$) of UMA-S-1P1 among different heads based on the MAD test dataset.

### v2

@@ III.2 Comparing model variants
Beyond the comparison between entirely different uMLIPs, we assess the consistency of information content across uMLIP “variants” of the same architecture. Through this analysis, we establish the sensitivity of various architectures to changes in the training dataset or target. We also establish a scale for the “internal” variability of each architecture that helps calibrate the comparison between models that differ more substantially. We categorize the variants into: “single-task” models where an entirely separate model architecture is trained from scratch for distinct datasets, and “multi-task” models that either adopt a multi-head Caruana (1997); Ruder (2017) or mixture-of-experts Jacobs et al. (1991); Riquelme et al. (2021); Shazeer et al. (2017) approach to simultaneously target multiple datasets with a single model architecture, often with significant weight sharing between the prediction tasks.
@@
Figure 2: The reconstruction errors across single-task MACE models trained on MPtrj (MACE-MP-0b3), OMat24, and MatPES (PBE and r2SCAN) datasets Deng et al. (2023); Kaplan et al. (2025); Barroso-Luque et al. (2024) evaluated on the MAD test subset.
@@
We first compare the reconstruction errors across the single-task MACE variants, which have been trained separately on MPtrj Deng et al. (2023), OMat24 Barroso-Luque et al. (2024), MatPES-PBE, and MatPES-r2SCAN Kaplan et al. (2025). The latter two datasets only differ in the exchange-correlation functional used during the reference calculations, which are PBE Perdew et al. (1996) and r2SCAN Furness et al. (2020), respectively. Overall, the reconstruction errors are moderately low across the variants (Fig. 2) (the average off-diagonal GFRE is 0.31, and the average LFRE is 0.12). The MPtrj variant exhibits the largest overall GFREs and LFREs (average GFRE and LFRE are 0.42 and 0.15, respectively), which reflects the limited information content of the smallest and oldest MPtrj dataset. The reconstruction errors between MatPES-PBE and MatPES-r2SCAN are comparable to the other cases, revealing that having different functionals in the reference calculations converges the resulting variant models into significantly distinct latent spaces. The lowest overall errors are observed for the OMat24 variant (on average, the GFRE is 0.25 and the LFRE is 0.09), which is indicative of the relatively rich information content of OMat24 dataset that is often utilized in the pre-training of uMLIPs Fu et al. (2025); Lysogorskiy et al. (2025); Batatia et al. (2025b); Rhodes et al. (2025); Barroso-Luque et al. (2024).
@@
In DPA-3.1, a model trained on the OpenLAM-v1 collection of 31 datasets Peng et al. (2025b), the following multi-task learning strategy is adopted: the model incorporates a shared backbone network while still being conditioned for different datasets via one-hot encoding. This backbone is followed by multiple specialized output heads, or branches, tailored to each dataset Zhang et al. (2025). Reconstruction results between branches corresponding to OMat24, MPtrj, OC20, ODAC, and SPICE datasets Barroso-Luque et al. (2024); Deng et al. (2023); Chanussot et al. (2021); Sriram et al. (2023); Eastman et al. (2023), presented in SI VIII.1.1, indicate comparable latent space consistency to the MACE single-task models, suggesting that strong parameter sharing without dynamic routing preserves highly universal latent representations.
@@
UMA is a uMLIP developed with the goal of having a single model that performs well across all atomistic modeling tasks, encompassing materials, molecules, and their interactions. To do so, the model has been trained simultaneously on OMol25, OMat24, OC20, ODAC25, and OMC25 Levine et al. (2025); Barroso-Luque et al. (2024); Chanussot et al. (2021); Sriram et al. (2023); Gharakhanyan et al. (2025). Instead of adopting a multi-head approach for the different datasets, the model incorporates a MoLE strategy in its interaction layers, where the weights of the different experts are computed based on the global embedding that encodes the dataset-specific information Wood et al. (2025). UMA-S-1P1 shows greater variability in reconstruction errors (see Fig. 3), which suggests that the MoLE mechanism encourages stronger specialization of feature manifolds corresponding to each dataset. We observe that the OMat24 task reconstructs other variants with relatively low errors (average off-diagonal GFRE is 0.49 and LFRE is 0.34), similar to previous results with MACE, whereas other variants struggle to approximate its latent space linearly. We also observe group-level consistency, as the features of catalysis-related tasks (OC20, ODAC25, OMC25) generally exhibit lower reconstruction errors (GFRE is 0.46, and LFRE is 0.34 on average across this group).
@@
Figure 3: The reconstruction errors across different input tasks of UMA-S-1P1 trained on OMat24, OMOL, OC20, ODAC, and OMC Barroso-Luque et al. (2024); Levine et al. (2025); Chanussot et al. (2021); Sriram et al. (2023); Gharakhanyan et al. (2025) evaluated on the MAD test subset.
@@
To investigate the influence of the training domain on the learned latent space, we also compare the organics-focused MACE-OFF23 model Kovács et al. (2025) with the materials-focused MACE-MP-0b3 Batatia et al. (2025a) by evaluating the reconstruction errors on the organic structures of the MAD test subset. As detailed in SI VIII.2 (Table S2), the materials-trained model reconstructs the organic-trained model almost perfectly (GFRE is 0.02), whereas the reverse direction yields remarkably higher global error (GFRE is 0.36), despite similar local agreement (LFRE is 0.04). This could suggest that the broader materials manifold learned by MACE-MP-0b3 cannot be fully captured by MACE-OFF23, which has been exposed only to organic systems.
@@
We further examine the effect of model capacity by comparing the MACE checkpoints of varying sizes (small, medium, large) within the same architecture family, for both the materials-focused MACE-MP-0a series and the organic-focused MACE-OFF23 series (SI VIII.3, Fig. S6). Interestingly, regardless of size, all MACE-MP-0a variants reconstruct each other with near-perfect linear mapping, whereas reconstructing the large MACE-OFF23 checkpoint from its smaller counterparts is significantly harder (average GFRE is 0.48 when small/medium $\rightarrow$ large). A similar size-scaling analysis on PET-OMAT models Laboratory of Computational Science and Modeling (COSMO), EPFL (2026) (SI Fig. S7) shows that smaller models can still approximate the principal manifold of larger ones reasonably well, with the medium-sized variant offering a good information content/size tradeoff.
@@
Finally, we assessed the impact of single-task versus multi-head readout architectures using two MACE models trained on the same OMat24 dataset: the original single-head MACE-OMAT-0 and the OMat-specific head of the multi-head MACE-MH-1 model Batatia et al. (2025b) (the latter having undergone multi-head replay fine-tuning after initial pre-training). Results are presented in SI VIII.4 (Table S4). Both directions exhibit moderate global reconstruction error (GFRE is $\sim$ 0.33) but low local error (LFRE is less than 0.10), demonstrating that the multi-head replay protocol preserves the core local feature neighborhoods of the original single-head representations while allowing the readout to become more expressive.
@@ III.3 Varying targets
Although we have primarily focused on uMLIPs, universal models are also available for properties other than the ground state potential energy surface. To this end, we compare the latent features of PET-MAD and PET-MAD-DOS How et al. (2025b). The two models have nearly identical architectures, differing only in the last linear readout layer due to the different sizes of the targets (last-layer features are still of the same dimension), and are both trained on the MAD dataset. While PET-MAD was trained on energies and forces, PET-MAD-DOS was trained on the electronic density of states (DOS).
The results in SI IX reveal that PET-MAD-DOS exhibits more information content than PET-MAD, as seen by the lower GFRE and LFRE values (0.68 and 0.39 for energy $\to$ DOS vs. 0.56 and 0.28 for DOS $\to$ energy). While this observation is expected, provided that DOS contains far more information than the energies and forces, it is still interesting to see that the information content in the latent features also reflects our physical knowledge of the targets. Still, the substantial reconstruction error in the DOS $\to$ energy direction indicates that the energy-based latent features also encode distinct information.
@@ III.4 Fine-tuning
The advent of uMLIPs has made fine-tuning on domain-specific datasets a standard approach for obtaining accurate potentials tailored to specific applications that can still generalize well Deng et al. (2024); Jacobs et al. (2025); Radova et al. (2025); Kaur et al. (2025). Here, we analyze the reconstruction errors between features that arise from different fine-tuning strategies, as well as a model trained entirely from scratch. We also study the evolution of learned latent features along a “training trajectory” by computing the reconstruction errors of latent features sampled from model checkpoints at intermediate training epochs with respect to the fully trained model.
@@
For our analysis, we consider the PET architecture and its pre-trained PET-MAD uMLIP as the baseline. We train material-specific MLIPs for lithium thiophosphate (LPS), a promising class of solid-state electrolytes widely studied for the next generation of battery materials Liu et al. (2013); Fragapane and Deringer (2025); Kato et al. (2021); Manthiram et al. (2017); Tisi et al. (2024); Gigli et al. (2024). We consider the following training or fine-tuning procedures: a bespoke model trained from scratch on the dataset from randomly initialized weights; full fine-tuning (FF), where all trainable model weights are optimized on the new dataset; “head-only” fine-tuning (HF), where only the weights of the MLP head are optimized during training while the rest remains fixed; full transfer learning (FTL), where a new MLP head is added with random weight initialization, and then all trainable weights of the model are optimized; head transfer learning (HTL), where a new MLP head is created with random weights, and only the MLP head weights are optimized. A schematic overview of the trainable parts for these fine-tuning strategies is given in Fig. 4. The training procedures were carried out using metatrain Bigi et al. (2025) (see SI X for the error metrics and hyperparameters of the fine-tuned models). We use the dataset from Ref. Türk et al. (2025), which was previously used to model complex surface structural reconstructions on LPS. The dataset contains 4,088 structures, yielding 225,905 atomic environments in total. In particular, it contains the structures of bulk LPS in each of its phases ($\gamma$, $\beta$, $\alpha$, and amorphous) and the structures of the LPS surface with and without absorbed water molecules.
@@
Figure 4: A schematic overview of the fine-tuning strategies for the PET architecture: full fine-tuning, head fine-tuning, full transfer learning, and head transfer learning. The graph neural network (GNN) based on the Point Edge Transformer (PET) computes the backbone (BB) features. For a given prediction target, these BB features are fed into the readout “heads” (multi-layer perceptrons or MLPs) to generate the last-layer (LL) features. The four illustrated strategies involve training different combinations of these model components, and the trainable parts are colored in salmon.
@@
Fig. 5 shows the reconstruction errors among the final models from the described fine-tuning strategies, the pre-trained PET-MAD model, and a bespoke PET model trained from scratch on the LPS dataset. The reconstruction errors are notably low between all of the fine-tuned models (FF, HF, FTL, and HTL) and the pre-trained PET-MAD model, which signifies a high degree of similarity in their last-layer features despite differences in training protocols. This suggests that the fine-tuned models retain a strong bias from the pre-trained model, which is also evident in the PCovR projections of the last-layer features (see Fig. S8). The bespoke model exhibits the highest reconstruction errors, yet the values are still moderate, revealing that its training has converged to a distinct yet nearby minima with respect to the other models.
@@
Figure 5: The reconstruction errors across the last-layer atom-level features of differently fine-tuned PET checkpoints computed for the LPS dataset.
@@
Fig. 6 shows the evolution of the reconstruction errors along the training trajectory computed relative to the fully converged model for each strategy. The bespoke model trained from scratch starts with high GFRE and LFRE, which gradually decay, yet remain above numerical zero even after 1000 epochs (0.009 for GFRE and 0.005 for LFRE). In contrast, fine-tuned strategies begin with lower initial reconstruction errors due to pre-conditioning with the PET-MAD uMLIP, which then converge quickly to near-zero values. In FTL and HTL, the curve is not monotonic, and more than 400 epochs are needed for convergence, most likely due to the randomized head initialization. In HF, a similar number of epochs is needed to align the features with those of the final model, whereas FF converges the reconstruction errors to <0.01<0.01 with only 20 epochs. This can be explained by the relative differences in the training degrees of freedom between the two strategies. Overall, the combination of low reconstruction errors in Fig. 5 and rapid convergence in Fig. 6 demonstrates that the last-layer features of PET-MAD provide a robust starting point for the fine-tuning strategies. This provides further explanation as to why the uMLIPs can converge rapidly to high accuracy when fine-tuned to a new, narrower dataset Kaur et al. (2025).
@@
Figure 6: The evolution of the reconstruction errors of last-layer atom-level features during training on the LPS dataset, where the errors have been computed with respect to the features of fully trained models corresponding to each respective fine-tuning strategy as targets.
@@
While we have chosen the last-layer (LL) features for our analysis thus far, the latent features can be extracted from any other layer within the NN architectures of the uMLIPs. Another features of particular interest are the “backbone” (BB) latent features, which correspond to the representations obtained immediately after all message-passing iterations, but before entering a multilayer perceptron (MLP) head, which is an architectural choice followed by many uMLIPs (see Fig. 4 for a sample schematic of PET-MAD). The BB features are of particular interest in multi-head and multi-task models, where they often serve as a shared input across the multiple heads. To provide examples of both multi-task and single-task contexts, we evaluate the reconstruction errors between the last-layer features and the backbone features for UMA and PET.
@@
For PET (Fig. S9 in SI XI), both reconstruction errors are higher and more varied across model variants compared to UMA-S-1P1, indicating greater divergence between BB and LL representations. Even in the GFRE, reconstructing BB from LL yields higher errors than the reverse case. Once again, the LFRE is lower in both reconstruction tasks, but the improvement is only minor in the case of LL $\rightarrow$ BB. We speculate that the large information loss might be associated with the MLP eliminating the non-invariant parts of the BB features, which takes place as the PET model learns the symmetries of the energy prediction task.
@@
Figure 7: The global and local reconstruction errors between the last-layer features ($L$) and backbone features ($B$) of UMA-S-1P1 for different experts, computed for the MAD test subset.